If an employee deposits Rs. 2,000 at the end of each year into his company’s plan which pays 7% interest compounded quarterly, how much will he have in the account at the end of 5 years?
1
Expert's answer
2019-05-21T09:37:29-0400

The formula for compound interest


"A=P(1+{r \over n})^{nt}"

A= the future value of the investment/loan, including interest

P = the principal investment amount

r = the annual interest rate (decimal)

n = the number of times that interest is compounded per unit t

t = the time the money is invested or borrowed for

We have that "P_0=2000, r=0.07, n=4"


"A_2=P_0\Big(1+(1+{r \over n})^n \Big)""A_3=P_0\Big(1+(1+{r \over n})^n+(1+{r \over n})^{2n} \Big)"

We see the geometric series


"S_k=1\cdot{1-\Big((1+{r \over n})^n \Big) ^k\over 1-(1+{r \over n})^n}"

Then


"A_m=P_0S_m=P_0\cdot{1-\Big((1+{r \over n})^n \Big) ^m\over 1-(1+{r \over n})^n}"

How much will he have in the account at the end of 5 years?


"A_6=2000\cdot{1-\Big((1+{0.07 \over 4})^4 \Big) ^6\over 1-(1+{0.07 \over 4})^4}=14373.78"

He will have in the account at the end of 5 years RM 14373.78.